Two - Dimensional Motion: Projectile

1. An object is launched at a velocity of 20 m/s in a direction making an angle of 25 upward with the horizontal.

A. What is the maximum height reached by the object?
B. What is the total fight time (between launch and touching the ground) of the object?
C. What is the magnitude of the velocity of the object just before it hits the ground?
1
Expert's answer
2021-09-26T08:40:50-0400

A. The height is


"h=\frac{(v\sin\theta)^2}{2g}=3.65\text{ m}."

B. The time is


"T=\frac{2v\sin\theta}{g}=1.72\text{ s}."

C. The velocity is almost the same, 20 m/s in magnitude, but 25° below the horizontal.